# The inertia bound is far from tight

### v2

@@ 1 Introduction
The ratio bound was first proved by Hoffman (unpublished), who only stated it in the case that $A$ is the (ordinary) adjacency matrix of $G$. Even this result is surprisingly powerful; for example, it can be used to give a short proof of the Erdős–Ko–Rado theorem [21] (see e.g. [23, Section 2.5] for details). However, the flexibility of choosing an arbitrary weighted adjacency matrix makes the ratio bound much more powerful (see e.g. [47]). Determining the optimal weights to use for the ratio bound can be formulated as a semidefinite program, and the rich theory of semidefinite optimization can therefore be used to theoretically and computationally determine the optimal choice of weights for any given graph.
@@
In many ways, the story of the inertia bound parallels that of the ratio bound. It was first proved by Cvetković [12], who stated it only for adjacency matrices. This already has a number of interesting applications, including another short proof of the Erdős–Ko–Rado theorem (see e.g. [23, Section 2.10] for details), but the flexibility of general weighted adjacency matrices makes the bound much more powerful. The general statement for arbitrary weighted adjacency matrices was perhaps first noted by Calderbank and Frankl [9], who used it to prove several results in extremal set theory.
@@
Sinkovic’s proof involves a great deal of casework, and as such is highly specific to the 17-vertex Paley graph. However, in recent years, new techniques were developed which shed more light on Godsil’s question. First, Mančinska and Roberson [39] introduced a new graph parameter, the quantum independence number $\alpha_{q}$, which satisfies $\alpha_{q}(G)\geq\alpha(G)$ for every graph $G$. Moreover, they proved [39, Corollary 4.8] that for infinitely many $n$, there exists an $n$-vertex graph with independence number $O(n^{1-\varepsilon})$ and quantum independence number $\Omega(n/{\log n})$, for some small absolute constant $\varepsilon>0$. More recently, Wocjan, Elphick and Abiad [48] proved that the inertia bound is also an upper bound on the quantum independence number, even when one allows Hermitian weights. Namely, they proved [48, Theorem 3.3] that if $A$ is a Hermitian weighted adjacency matrix [^2] Here, we say that $A$ is a Hermitian weighted adjacency matrix of $G$ if it is a Hermitian matrix with complex entries such that $A_{ij}=0$ whenever $ij\notin E(G)$. of $G$, then $\alpha_{q}(G)\leq n_{\geq 0}(A)$. Putting these results together, one obtains the following, stronger, negative answer to 1.3.
@@
Third, even though we now know that the inertia bound is not always tight, perhaps it is still the case that the inertia bound is in some sense “the best possible spectral bound” on the independence number. Concretely, is the inertia bound always stronger than the ratio bound? Actually, this question is closely related to the previous one, since the ratio bound is also an upper bound on the quantum independence number of a graph [39, Corollary 4.9].
@@
In fact, our main technical result is substantially stronger than Theorem 1.6: it shows that the inertia bound is very weak whenever $G$ is $C_{4}$-free, i.e. does not contain $C_{4}$ as a subgraph. We can moreover obtain a small constant-factor improvement when $G$ is also $C_{3}$-free, i.e. has girth at least 5.
@@
Theorem 1.7 immediately implies the existence of graphs for which the inertia bound is far from tight, by letting $G$ be a $C_{4}$-free graph whose independence number is $o(n)$. There are many examples of such graphs, defined explicitly or obtained via the probabilistic method. In particular, the statement of Theorem 1.6 follows from 1.7(ii), taking $G$ to be an edge-transitive induced subgraph of at least half the vertices of the polarity graph of a finite projective plane, with girth at least 5 (such a graph was shown to exist by Parsons [42, Theorem 1(a)]) [^4] By [37, Theorem 9], the Lovász theta number of $G$ is then $|\lambda_{\mathrm{min}}/(\lambda_{\mathrm{max}}-\lambda_{\mathrm{min}})|n$, where $\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}$ are the minimum and maximum eigenvalues, respectively, of the ordinary adjacency matrix. The Lovász theta number is monotonic with respect to induced subgraphs, and it is well-known (e.g. [41]) that the theta number of full polarity graph $G_{0}$ is at most $2\lvert V(G_{0})\rvert^{3/4}$.. It is an interesting byproduct of our results that the inertia bound is so weak for such algebraically-defined graphs, as generally speaking, techniques from spectral graph theory seem particularly well-suited to highly structured graphs with many symmetries.
@@ 2.1 Probabilistic bounds via moments
One can get stronger bounds if one is given information on $\mathbb{E}[Y^{3}]$; the following such result is a special case of [49, equation (12)], specialized to $\mathbb{E}[Y^{3}]=0$.
Lemma 2.2 ([49, equation (12)]).
@@ 2.2 Simple facts about eigenvalues
We will need two basic results of linear algebra. The first is called Sylvester’s law of inertia; see e.g. [27, Theorem 4.5.8] for a proof.
@@
We will also use the following simple consequence of Cauchy’s interlacing formula; see e.g. [24, Theorem 9.1.1] or [27, Theorem 4.3.28] for a proof. We remark that it immediately implies the inertia bound if applied to a zero principal submatrix, i.e. an independent set.
@@ 2.3 Matrix scaling
Now, let $H$ be the subgraph of $G$ induced on the vertex set $(A\setminus\{a\})\cup(B\setminus\{b\})$. Then $H$ has no perfect matching. By Hall’s theorem (see e.g. [24, Theorem 15.15.3]), this implies that there exists some non-empty $S\subseteq A$ with $\lvert N_{H}(S)\rvert<\lvert S\rvert$. Letting $T=(B\setminus\{b\})\setminus N_{H}(S)$, we find that there is no edge between $S$ and $T$, and
@@
where $A_{1},A_{2}$ are principal submatrices of $A$ of sizes $\lvert S\setminus T\rvert,\lvert T\setminus S\rvert$, respectively, and where *** denotes an arbitrary submatrix. By Lemma 2.5, we know that $n_{\geq 0}(A)\geq n_{\geq 0}(A^{\prime})$, where $A^{\prime}$ is the principal submatrix on rows and columns in $S\cup T$. Moreover, as $A^{\prime}$ is block-diagonal, its eigenvalues are those of $A_{1}$, those of $A_{2}$, plus $\lvert S\cap T\rvert$ additional zero eigenvalues. Applying the inductive hypothesis to $A_{1}$ and $A_{2}$, which are Hermitian weighted adjacency matrices of smaller $C_{4}$-free graphs, we find that
@@ 4 Concluding remarks
For any $\varepsilon>0$, the following holds with probability $1-o(1)$ for $G\sim\mathbb{G}(n,\frac{1}{2})$ as $n\to\infty$. Every weighted adjacency matrix $A$ of $G$ has $n_{\geq 0}(A)\geq(\frac{1}{2}-\varepsilon)n$.
In particular, as $\alpha(\mathbb{G}(n,\frac{1}{2}))=O(\log n)$ with probability $1-o(1)$, 4.1 would imply that the inertia bound is far from tight for almost all graphs.
@@
We remark that a result of Konyagin [34] implies the analogous statement for the ratio bound, i.e. that for every $k\in\mathbb{N}$ there exists a graph $G$ with $\alpha(G)=2$ for which the ratio bound (with any weighted adjacency matrix) cannot be used to prove $\alpha(G)<k$. Indeed, Konaygin showed that there are graphs $G$ with $\alpha(G)=2$ which have arbitrarily large values of the so-called Lovász theta number [37]. For more details, see [1] or [38, Chapter 11].
@@ Acknowledgements:
We are grateful to Anurag Bishnoi, Clive Elphick, and Ferdinand Ihringer for helpful comments and interesting discussions on an earlier draft of this paper.

### v3

@@ 1 Introduction
The ratio bound was first proved by Hoffman (unpublished), who only stated it in the case that $A$ is the (ordinary) adjacency matrix of $G$. Even this result is surprisingly powerful; for example, it can be used to give a short proof of the Erdős–Ko–Rado theorem [22] (see e.g. [24, Section 2.5] for details). However, the flexibility of choosing an arbitrary weighted adjacency matrix makes the ratio bound much more powerful (see e.g. [48]). Determining the optimal weights to use for the ratio bound can be formulated as a semidefinite program, and the rich theory of semidefinite optimization can therefore be used to theoretically and computationally determine the optimal choice of weights for any given graph.
@@
In many ways, the story of the inertia bound parallels that of the ratio bound. It was first proved by Cvetković [13], who stated it only for adjacency matrices. This already has a number of interesting applications, including another short proof of the Erdős–Ko–Rado theorem (see e.g. [24, Section 2.10] for details), but the flexibility of general weighted adjacency matrices makes the bound much more powerful. The general statement for arbitrary weighted adjacency matrices was perhaps first noted by Calderbank and Frankl [10], who used it to prove several results in extremal set theory.
@@
Sinkovic’s proof involves a great deal of casework, and as such is highly specific to the 17-vertex Paley graph. However, in recent years, new techniques were developed which shed more light on Godsil’s question. First, Mančinska and Roberson [40] introduced a new graph parameter, the quantum independence number $\alpha_{q}$, which satisfies $\alpha_{q}(G)\geq\alpha(G)$ for every graph $G$. Moreover, they proved [40, Corollary 4.8] that for infinitely many $n$, there exists an $n$-vertex graph with independence number $O(n^{1-\varepsilon})$ and quantum independence number $\Omega(n/{\log n})$, for some small absolute constant $\varepsilon>0$. More recently, Wocjan, Elphick and Abiad [49] proved that the inertia bound is also an upper bound on the quantum independence number, even when one allows Hermitian weights. Namely, they proved [49, Theorem 3.3] that if $A$ is a Hermitian weighted adjacency matrix [^2] Here, we say that $A$ is a Hermitian weighted adjacency matrix of $G$ if it is a Hermitian matrix with complex entries such that $A_{ij}=0$ whenever $ij\notin E(G)$. of $G$, then $\alpha_{q}(G)\leq n_{\geq 0}(A)$. Putting these results together, one obtains the following, stronger, negative answer to 1.3.
@@
Third, even though we now know that the inertia bound is not always tight, perhaps it is still the case that the inertia bound is in some sense “the best possible spectral bound” on the independence number. Concretely, is the inertia bound always stronger than the ratio bound? Actually, this question is closely related to the previous one, since the ratio bound is also an upper bound on the quantum independence number of a graph [40, Corollary 4.9].
@@
In fact, our main technical result is substantially stronger than Theorem 1.6: it shows that the inertia bound is very weak whenever $G$ is $C_{4}$-free, i.e. does not contain $C_{4}$ as a subgraph. Moreover, we can obtain a small constant-factor improvement when $G$ is also $C_{3}$-free, i.e. has girth at least 5.
@@
Theorem 1.7 immediately implies the existence of graphs for which the inertia bound is far from tight, by letting $G$ be a $C_{4}$-free graph whose independence number is $o(n)$. There are many examples of such graphs, defined explicitly or obtained via the probabilistic method. In particular, the statement of Theorem 1.6 follows from 1.7(ii), taking $G$ to be an edge-transitive induced subgraph of at least half the vertices of the polarity graph of a finite projective plane, with girth at least 5 (such a graph was shown to exist by Parsons [43, Theorem 1(a)]) [^4] By [38, Theorem 9], the Lovász theta number of $G$ is then $|\lambda_{\mathrm{min}}/(\lambda_{\mathrm{max}}-\lambda_{\mathrm{min}})|n$, where $\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}$ are the minimum and maximum eigenvalues, respectively, of the ordinary adjacency matrix. The Lovász theta number is monotonic with respect to induced subgraphs, and it is well-known (e.g. [42]) that the theta number of the full polarity graph $G_{0}$ is at most $2\lvert V(G_{0})\rvert^{3/4}$.. It is an interesting byproduct of our results that the inertia bound is so weak for such algebraically-defined graphs, as generally speaking, techniques from spectral graph theory seem particularly well-suited to highly structured graphs with many symmetries.
@@ 2.1 Probabilistic bounds via moments
One can get stronger bounds if one is given information on $\mathbb{E}[Y^{3}]$; the following such result is a special case of [50, equation (12)], specialized to $\mathbb{E}[Y^{3}]=0$.
Lemma 2.2 ([50, equation (12)]).
@@ 2.2 Simple facts about eigenvalues
We will need two basic results of linear algebra. The first is called Sylvester’s law of inertia; see e.g. [28, Theorem 4.5.8] for a proof.
@@
We will also use the following simple consequence of Cauchy’s interlacing formula; see e.g. [25, Theorem 9.1.1] or [28, Theorem 4.3.28] for a proof. We remark that it immediately implies the inertia bound if applied to a zero principal submatrix, i.e. an independent set.
@@ 2.3 Matrix scaling
Now, let $H$ be the subgraph of $G$ induced on the vertex set $(A\setminus\{a\})\cup(B\setminus\{b\})$. Then $H$ has no perfect matching. By Hall’s theorem (see e.g. [25, Theorem 15.15.3]), this implies that there exists some non-empty $S\subseteq A$ with $\lvert N_{H}(S)\rvert<\lvert S\rvert$. Letting $T=(B\setminus\{b\})\setminus N_{H}(S)$, we find that there is no edge between $S$ and $T$, and
@@
where $A_{1},A_{2}$ are principal submatrices of $A$ of sizes $\lvert S\setminus T\rvert,\lvert T\setminus S\rvert$, respectively, and where * denotes an arbitrary submatrix. By Lemma 2.5, we know that $n_{\geq 0}(A)\geq n_{\geq 0}(A^{\prime})$, where $A^{\prime}$ is the principal submatrix on rows and columns in $S\cup T$. Moreover, as $A^{\prime}$ is block-diagonal, its eigenvalues are those of $A_{1}$, those of $A_{2}$, plus $\lvert S\cap T\rvert$ additional zero eigenvalues. Applying the inductive hypothesis to $A_{1}$ and $A_{2}$, which are Hermitian weighted adjacency matrices of smaller $C_{4}$-free graphs, we find that
@@ 4 Concluding remarks
The following holds with probability $1-o(1)$ for $G\sim\mathbb{G}(n,\frac{1}{2})$ as $n\to\infty$. Every weighted adjacency matrix $A$ of $G$ has $n_{\geq 0}(A)=\Omega(\frac{n}{\log n})$.
In particular, as $\alpha(\mathbb{G}(n,\frac{1}{2}))=O(\log n)$ with probability $1-o(1)$, 4.1 would imply that the inertia bound is far from tight for almost all graphs. An earlier version of this paper conjectured a stronger bound, namely that $n_{\geq 0}(A)\geq(\frac{1}{2}-o(1))n$ for every weighted adjacency matrix of $G\sim\mathbb{G}(n,\frac{1}{2})$. However, Noga Alon pointed out to us that such a statement is false, and that 4.1 would be best possible if true. Alon observed that for every graph $G$, there exists a weighted adjacency matrix $A$ with $n_{\geq 0}(A)=\chi(\overline{G})$, where $\chi(\overline{G})$ is the clique cover number of $G$. Indeed, one can easily check that the adjacency matrix of a disjoint union of $m$ cliques has exactly $m$ non-negative eigenvalues, which implies the claim by taking $A$ to be the adjacency matrix of a subgraph of $G$ which forms an optimal clique cover. A famous result of Bollobás [7] implies that $\chi(\overline{G})=O(\frac{n}{\log n})$ with probability $1-o(1)$ for $G\sim\mathbb{G}(n,\frac{1}{2})$, showing that 4.1 would be best possible if true.
@@
We remark that a result of Konyagin [35] implies the analogous statement for the ratio bound, i.e. that for every $k\in\mathbb{N}$ there exists a graph $G$ with $\alpha(G)=2$ for which the ratio bound (with any weighted adjacency matrix) cannot be used to prove $\alpha(G)<k$. Indeed, Konaygin showed that there are graphs $G$ with $\alpha(G)=2$ which have arbitrarily large values of the so-called Lovász theta number [38]. For more details, see [1] or [39, Chapter 11].
@@ Acknowledgements:
We are grateful to Noga Alon, Anurag Bishnoi, Clive Elphick, and Ferdinand Ihringer for helpful comments and interesting discussions on earlier drafts of this paper.